OpenStudy (clamin):

PLEASE HELP!! MEDAL!! SOLVE FOR X:

OpenStudy (clamin):

\[\sqrt[3]{4X+1}=5\]

OpenStudy (theedwardsfamily):

http://www.symbolab.com/

OpenStudy (nincompoop):

undo the root

OpenStudy (nincompoop):

get the cube of each side so it undoes the root on the left-hand side and then try to solve for x

OpenStudy (clamin):

@TheEdwardsFamily I cant write 3 inside the square root

OpenStudy (theedwardsfamily):

oh

OpenStudy (clamin):

what do you get the cube in each side?? @nincompoop

OpenStudy (clamin):

a) x=-31 b)x=6 c)x=31 d)no real solution

OpenStudy (clamin):

i think its d

OpenStudy (nincompoop):

did you know that your problem is the same as \( (4x+1)^{\frac{1}{3}}=5^\frac{1}{1}\) if you cube both sides, you are essentially \( (4x+1)^{\frac{1}{3} \times \frac{3}{1}}=5^{\frac{3}{1}}\)

OpenStudy (clamin):

@Nnesha

OpenStudy (nincompoop):

@sleepyjess show your skills

OpenStudy (sleepyjess):

huh? why me? :(

Nnesha (nnesha):

yes @claim ??

OpenStudy (nincompoop):

try to pick up where I left off, @sleepyjess

OpenStudy (clamin):

i have no clue what is this

OpenStudy (nincompoop):

I ALREADY GAVE YOU THE CLUE!

Nnesha (nnesha):

okay so you have to take cube both side so you can cancel cube root on left side

OpenStudy (nincompoop):

do you not know how to multiply, add, subtract and divide? learn the laws of exponents!

OpenStudy (sleepyjess):

Ok, so nin is basically saying that your problem is the left side to the power of 1/3 and the right side is just to the power of 1. Now cube both sides to get ride of 1/3 and make it 1.

OpenStudy (clamin):